# MSTAR: Box-free Multi-query Scene Text Retrieval with Attention Recycling

### v1

@@ 4.2 Implementation details
The visual encoder $\phi$ is initialized from ViT-Base-512 of SigLIP [49]. The multi-modal encoder $\psi$ is initialized from BLIP2 [19]. The number of query tokens $Q_{\text{l}}$ is set to 64 with interpolation, which is consistent with the setting of the vanilla BLIP2 in our comparison experiments. The weights of the MLP, SAS, and MIM modules are randomly initialized. The MSTAR model was trained on four NVIDIA A800 GPUs and evaluated on a single GPU, using the AdamW optimizer [25]. A multi-stage training is adopted with progressive resolution increasing from 512×512, 640×640, to 800×800. For re-ranking, the top 2% of images are selected from the initial retrieval results. For instance, in a dataset containing $10k$ images, only the top 200 images are used for re-ranking.
@@
Total-Text [7] IJDAR’20 60 300
CTW [21] PR’19 100 500
IC15 [14] ICDAR’15 100 500
CTR [36] Arxiv’16 500 7196
STR [9] ECCV’18 50 10k
CSVTR [37] CVPR’21 23 1667
PSTR [48] ACM MM’24 36 1080
@@
ABCNet [22] TPAMI’21 24.13 26.14 15.15 36.47 18.74
MaskTextSpotter [20] ECCV’20 32.43 46.72 27.53 29.08 26.37
TDSL [37] CVPR’21 58.25 69.11 40.83 72.71 50.36
Deepsolo [46] CVPR’23 52.04 67.54 25.68 72.14 42.79
TG-Bridge [11] CVPR’24 54.09 69.89 30.21 75.53 40.73
@@
SPTSv2 [23] TPAMI’23 35.18 33.56 21.24 50.76 35.16
BLIP2  [19] PMLR’23 36.13 17.31 32.76 25.80 68.63
SigLIP  [49] CVPR’23 36.06 17.81 32.88 21.81 72.23
BLIP2 (FT) [19] PMLR’23 58.11 58.09 42.23 60.84 71.24
@@
Table 2: Evaluations of MAPon MQTR. FT denotes FineTune. The best results are shown in bold.
BLIP2 [19] TDSL [37] SigLIP [49] FDP [48] MSTAR
@@
Table 3: Comparisons on Prase-level Scene Text Retrieval dataset [48].
@@
Mishra et al. [27] ICCV’13 42.70 56.24 - - - - - 0.1
Jaderberg et al. [12] IJCV’16 86.30 66.50 - - - - - 0.3
Gomez et al. [9] ECCV’18 83.74 69.83 41.05 - - - - 43.5
Mafla et al. [26] PR’21 85.74 71.67 - - - - - 42.2
TDSL  [37] CVPR’21 89.38 77.09 66.45 74.75 59.34 77.67 74.16 12.0
Wang et al. [38] TPAMI’24 - 81.02 72.95 - - - - 9.3
Wen et al. [40] WSDM’23 90.95 77.40 - 80.09 - - - 11.0
FDP-RN50$\times$16 [48] ACM MM’24 89.63 89.46 - 79.18 - - - 11.8
@@
MSTAR (+rerank) - 91.11 86.14 65.25 86.96 92.95 82.69 84.18 6.9
Table 4: Comparisons with scene text retrieval methods of MAPon 6 public word retrieval datasets. The best results are shown in bold, and the second results are underlined.
@@
ABCNet [22] TPAMI’21 82.43 67.25 41.25 73.23 74.82 69.28 68.04 17.5
MaskTextspotterV3 [20] ECCV’20 83.14 74.48 55.54 83.29 80.03 77.00 75.58 2.4
Deepsolo [46] CVPR’23 87.15 76.58 67.22 83.19* 87.67* 82.80* 80.77 10.0
TG-Bridge [11] CVPR’24 87.23 81.30 70.08 87.11* 88.39* 83.55 * 82.94 6.7
@@
MSTAR (+rerank) - 91.11 86.14 65.25 86.96 92.95 82.69 84.18 6.9
Table 5: Comparisons with mainstream scene text spotting methods. * indicates that the model has been finetuned on the corresponding training set. The best results are highlighted in bold, and the second results are underlined.
@@ 4.3 Multi-Query Scene Text Retrieval
To enable multi-query retrieval, we collect a training dataset consisting of $95k$ images. First, $50k$ synthetic images with word transcriptions are leveraged from SynthText-900k [9]. Then $20k$ real images containing captions are collected from TextCap  [34]. We use the labels with both image captions and text transcriptions acquired with Rosetta  [3]. Besides, we have synthesized $25k$ images with phrase transcriptions using the synthesis engine [10]. Additionally, word or phrase annotations are utilized to form nonrepeated combined queries for images containing over one text instance. More training details can be found in the appendix.
On the MQTR dataset, we perform evaluation with both the representative box-based models and box-free models for multi-query scene text retrieval. For text spotting methods, we use the normalized edit distance to measure query-image matching scores following  [37]. To evaluate the models that can only handle word queries, we calculate the mean similarity of each word as the image-text scores for multi-word queries. The codes and weights are acquired from their official repositories.
Evaluation on multi-query scene text retrieval. As the results reported in Tab. 2, box-based methods [37, 46, 11] typically perform better on word queries and combined queries, which demand fine-grained scene text perception. However, these box-based models cannot leverage the rich linguistic semantics for phrase and semantic retrieval. Compared to box-based methods, MSTAR outperforms TG-Bridge[11] by 3.38% and Deepsolo[46] by 5.73% in word retrieval. On the other hand, VLMs such as BLIP2 (ViT-L) [19] and SigLIP (ViT-B-512) [49] perform better on phrase and semantic queries but struggle in word and combined retrieval. Compared to them, MSTAR outperforms SigLIP by 11.34% on phrase retrieval and 2.91% on semantic retrieval. On the overall results, our MSTAR obtains an improvement of 8.53% over previous works on average. These comparison results show the great advantages of our MSTAR on multi-query retrieval.
Additionally, MSTAR is also evaluated on the phrase-level scene text retrieval dataset  [48]. As shown in the Tab. 3, our MSTAR achieves 95.71% of MAP on the benchmark.
@@ 4.4 Word-level Scene Text Retrieval
In this part, we present comparison experiments on word-level retrieval. The model is trained on $100k$ images randomly sampled from SynthText-900k [9] and 5k images from MLT-5K [28] dataset. The evaluation setting keep the same as the previous method [37]. Note that Deepsolo  [46], TG-Bridge  [11], and SPTSv2 [23] are well finetuned on the training set of Total-Text, CTW and IC15 dataset.
Comparisons with text retrieval methods. We conduct comprehensive evaluations on the test sets of six public datasets, the results are presented in Tab. 4. Compared to FDP-RN50$\times$16 [48], our MSTAR achieves an improvement of 1.68% on SVT and 6.37% on Total-Text. While MSTAR demonstrates a slightly lower performance on the STR dataset, it eliminates the cost of expensive bounding-box for training. Compared to TDSL  [37], our method significantly outperforms the method across five datasets. Notably, our MSTAR surpasses TDSL by 9.16% on STR, 10.80% on TotalText, and 31.53% on CTW. These results verify the robust capabilities of our model. However, on the CTR dataset that contains extreme small text, our method underperforms TDSL. This is due to the absence of precise box supervision of our method, which is a common problem for box-free methods. Overall, MSTAR outperforms TDSL by an average 8.40% in MAP across six datasets. To further enhance performance, we rerank the top 2% of retrieved images by jointly feeding the text queries and images into the model. This re-ranking strategy improves performance by an additional 1.56% in MAP.
@@ 4.5 Ablation Study
Ablation study on three core components. The overall results are reported in Tab. 6. We begin by directly training the vision encoder, MLP and multi-modal encoder with standard contrastive learning. The results suggest that this baseline performs poorly on the fine-grained recognition capability, especially on the CTR and IC15 datasets which features small text. Then instructions are adopted to prompt the model to encode queries which leads to improvement. Subsequently, the MIM is added, which leads to a significant improvement of 2.19% on STR and 1.35% on CTW. However, a slight performance drop of 0.46% is observed on the MQTR dataset. This may be because of the use of Hungarian matching, which is effective for word-level queries and instance-level alignment, introduces confusion when handling more complex queries. Lastly, we incorporate PVE which is designed to improve the retrieval performance of the insalient objects such as small and detailed text instances. The results show a substantial improvement on CTR (4.36%) and IC15 (1.98%), validating the effectiveness of PVE in small text scenarios.
@@
TH+CC 59.66 85.17 80.17
TH+WS+CC 60.13 85.55 81.21
Table 7: Ablation studies on $\sigma$ algorithm. TH denotes ThresHolding, CC denotes Connected Components, and WS denotes WaterShed algorithm.
@@ 5 Discussion
Application of MSTAR for text localization. To further validate the effectiveness of MSTAR, we use it for text localization using Grad-CAM [32]. As illustrated in Fig.  3, MSTAR can localize different types of queries. For example, MSTAR accurately identifies the target text instance “dream big” on a t-shirt, distinguishing it from the “dream big” of the shorts in Fig.  3 (a). Besides, MSTAR can also accurately localize curved and dense text instances. As the Fig.  3 (e) shows, MSTAR successfully identifies the word “copyright” within a document page image. These results demonstrate that MSTAR can accurately localize text instances without box supervision.
Discussion about the pretrained model. We discuss how the pretrained model affects performance on the scale of parameters and data. We tested different variants of CLIP [43], BLIP [18], BLIP2 [19], and SigLIP [49] on the CTR dataset. The results show that despite the massive pretraining, the models struggle to achieve retrieval on small text instances. Details are provided in the appendix.
Limitations Although our method shows evident improvements in fine-grained scene text retrieval, it still faces challenges in handling extremely small and dense text instances. In particular, retrieval performance decreases substantially as the text size becomes smaller. However, this is a common limitation to box-free approaches.
@@ 6 Conclusion
In this paper, we introduce MSTAR, a new box-free method for multi-query scene text retrieval. It incorporates PVE to shift image attention to fine-grained scene text. Our model demonstrate competitive retrieval performance with state-of-the-art box-based methods while significantly reducing annotation cost. Moreover, for the first time, we study the multi-query scene text retrieval and establish the MQTR benchmark for evaluation. Experiments show neither box-based methods nor general cross-modal retrievers can handle such a challenging task while our MSTAR can serve as a strong baseline model for this task. For future work, we aim to investigate model acceleration and document retrieval tasks.
@@ A.1 Further Observations on VLMs
To validate the observations that VLMs tend to overlook detailed and insalient text instances, we conducted additional experiments with more models. For the BLIP, we calculate the ITM score which calculates the one-to-one matching probability between the image and the query. For the SigLIP models, we calculate the dot products between the vision and text embeddings without normalization following the official code. As illustrated in Fig. 4 (a), the BLIP can easily capture the salient text like “dead lakes park” but cannot find the text “old florida”. However, it can capture the “old florida” when the salient regions are covered. Similar observations are presented in Fig. 4 (b) (c).
@@ A.2 Performance of VLMs on Small Text Dataset
To test the pretrained ability of VLMs on finegrained perception, we evaluate representative VLMs (CLIP [43], BLIP [18], BLIP2 [19], SigLIP [49]) on the CTR [36] dataset. The CTR dataset is collected from the COCO dataset and features small scene text instances. As the results in Tab.  9, all of the tested VLMs struggle to retrieve images accurately. For instance, the CLIP variants obtain MAP of 8.1% while the BLIP gets 6.9% of MAP. Moreover, increasing model parameters and seen data does not lead to significant performance gains. Among these models, BLIP2 has the highest MAP of 13.3% of the tested models, indicating limited capacity of VLMs for small text retrieval.
Model Parameters Pretraining Data MAP%
@@
BLIP2-ViT-Large 452M 129M images 13.3
@@ B.1 Construction details of the MQTR dataset
Word Subset. The word subset includes 5000 images and 200 word queries. The images were sourced from the test set of SVT, CTW, IC15, Total-Text and the CTR. We extracted the word-level annotations from the datasets and filtered out the words with less than 3 characters (e.g. “st”). After that, 200 word queries were selected according to the word frequency.
Phrase Subset. The images of phrase subset include 1k images from PSTR [48], 1k images manually collected from the Web and 8k images from HierText [24]. Similarly, the line-level annotations were used and the lines with only one word were filtered out. Lastly, 200 phrase queries were selected according to frequency.
Combined Subset We first use all images collected from the CTR and HierText dataset. Given the word and line annotations of an image, we first filtered out the queries less than 3 characters. Then we implement an algorithm to compute all the combinations that contain 2-4 words. Top 200 text combinations on the dataset were selected according to frequency and repeat ratios. Then images paired to these queries were first selected as positive samples. Then images containing words similar to the 200 combined queries are also selected as negative samples according to the edit distance. Lastly, there are 5000 images for the positive samples and negative samplesin total.
Semantic Subset The semantic subset is manually collected from the web. we first brainstorm common used scene text queries and then search for candidate images from the web. In total, the subset consists of 25 queries and 1000 images.

### v2

@@ 4.2 Implementation details
The visual encoder $\phi$ is initialized from ViT-Base-512 of SigLIP [zhai2023sigmoid]. The multi-modal encoder $\psi$ is initialized from BLIP-2 [li2023blip]. The number of query tokens $Q_{\text{l}}$ is set to 64 with interpolation, which is consistent with the setting of the vanilla BLIP-2 in our comparison experiments. The weights of the MLP, SAS, and MIM modules are randomly initialized. The MSTAR model was trained on four NVIDIA A800 GPUs and evaluated on a single GPU, using the AdamW optimizer [loshchilov2017decoupled]. A multi-stage training is adopted with progressive resolution increasing from 512×512, 640×640, to 800×800. For re-ranking, the top 2% of images are selected from the initial retrieval results. For instance, in a dataset containing $10k$ images, only the top 200 images are used for re-ranking.
@@
Total-Text [ch2020total] IJDAR’20 60 300
CTW [liu2019curved] PR’19 100 500
IC15 [karatzas2015icdar] ICDAR’15 100 500
CTR [veit2016coco] Arxiv’16 500 7196
STR [gomez2018single] ECCV’18 50 10k
CSVTR [wang2021scene] CVPR’21 23 1667
PSTR [zeng2024focus] ACM MM’24 36 1080
@@
ABCNet [liu2021abcnet] TPAMI’21 24.13 26.14 15.15 36.47 18.74
MaskTextSpotter [liao2020mask] ECCV’20 32.43 46.72 27.53 29.08 26.37
TDSL [wang2021scene] CVPR’21 58.25 69.11 40.83 72.71 50.36
Deepsolo [ye2023deepsolo] CVPR’23 52.04 67.54 25.68 72.14 42.79
TG-Bridge [huang2024bridging] CVPR’24 54.09 69.89 30.21 75.53 40.73
@@
SPTSv2 [liu2023spts] TPAMI’23 35.18 33.56 21.24 50.76 35.16
BLIP-2  [li2023blip] PMLR’23 36.13 17.31 32.76 25.80 68.63
SigLIP  [zhai2023sigmoid] CVPR’23 36.06 17.81 32.88 21.81 72.23
BLIP-2 (FT) [li2023blip] PMLR’23 58.11 58.09 42.23 60.84 71.24
@@
Table 2: Evaluations of MAP% on MQTR. FT denotes fine-tune. The best results are shown in bold.
BLIP-2 [li2023blip] TDSL [wang2021scene] SigLIP [zhai2023sigmoid] FDP [zeng2024focus] MSTAR
@@
Table 3: Comparisons on Phrase-level Scene Text Retrieval dataset [zeng2024focus].
@@
Mishra et al. [mishra2013image] ICCV’13 42.70 56.24 - - - - - 0.1
Jaderberg et al. [jaderberg2016reading] IJCV’16 86.30 66.50 - - - - - 0.3
Gomez et al. [gomez2018single] ECCV’18 83.74 69.83 41.05 - - - - 43.5
Mafla et al. [mafla2021real] PR’21 85.74 71.67 - - - - - 42.2
TDSL  [wang2021scene] CVPR’21 89.38 77.09 66.45 74.75 59.34 77.67 74.16 12.0
Wang et al. [wang2024partial] TPAMI’24 - 81.02 72.95 - - - - 9.3
Wen et al. [wen2023visual] WSDM’23 90.95 77.40 - 80.09 - - - 11.0
[…]
@@
MSTAR (+re-rank) - 91.11 86.14 65.25 86.96 92.95 82.69 84.18 6.9
Table 4: Comparisons with scene text retrieval methods of MAP% on 6 public word retrieval datasets. The best results are shown in bold, and the second results are underlined.
@@
ABCNet [liu2021abcnet] TPAMI’21 82.43 67.25 41.25 73.23 74.82 69.28 68.04 17.5
MaskTextspotterV3 [liao2020mask] ECCV’20 83.14 74.48 55.54 83.29 80.03 77.00 75.58 2.4
Deepsolo [ye2023deepsolo] CVPR’23 87.15 76.58 67.22 83.19* 87.67* 82.80* 80.77 10.0
TG-Bridge [huang2024bridging] CVPR’24 87.23 81.30 70.08 87.11* 88.39* 83.55 * 82.94 6.7
@@
MSTAR (+re-rank) - 91.11 86.14 65.25 86.96 92.95 82.69 84.18 6.9
Table 5: Comparisons with mainstream scene text spotting methods. * indicates that the model has been fine-tuned on the corresponding training set. The best results are highlighted in bold, and the second results are underlined.
@@ 4.3 Multi-Query Scene Text Retrieval
To enable multi-query retrieval, we collect a training dataset consisting of $95k$ images. First, $50k$ synthetic images with word transcriptions are leveraged from SynthText-900k [gomez2018single]. Then $20k$ real images containing captions are collected from TextCap  [sidorov2020textcaps]. We use labels with both image captions and text transcriptions acquired with Rosetta  [borisyuk2018rosetta]. In addition, we have synthesized $25k$ images with phrase transcriptions using the synthesis engine [gupta2016synthetic]. Additionally, word or phrase annotations are utilized to form nonrepeated combined queries for images containing over one text instance. More training details can be found in the appendix.
On the MQTR dataset, we perform evaluation with both the representative box-based models and box-free models for multi-query scene text retrieval. For text spotting methods, we use the normalized edit distance to measure query-image matching scores following  [wang2021scene]. To evaluate the models that can only handle word queries, we calculate the mean similarity of each word as the image-text scores for multi-word queries. The codes and weights are acquired from their official repositories.
Evaluation on multi-query scene text retrieval. As the results reported in Tab. 2, box-based methods [wang2021scene, ye2023deepsolo, huang2024bridging] typically perform better on word queries and combined queries, which demand fine-grained scene text perception. However, these box-based models cannot leverage the rich linguistic semantics for phrase and semantic retrieval. Compared to box-based methods, MSTAR outperforms TG-Bridge[huang2024bridging] by 3.38% and Deepsolo[ye2023deepsolo] by 5.73% in word retrieval. On the other hand, VLMs such as BLIP-2 (ViT-L) [li2023blip] and SigLIP (ViT-B-512) [zhai2023sigmoid] perform better on phrase and semantic queries but struggle in word and combined retrieval. Compared to them, MSTAR outperforms SigLIP by 11.34% on phrase retrieval and 2.91% on semantic retrieval. In terms ofms of overall results, our MSTAR obtains an improvement of 8.53% over previous works on average. These comparison results show the great advantages of our MSTAR on multi-query retrieval.
Additionally, MSTAR is also evaluated on the phrase-level scene text retrieval dataset  [zeng2024focus]. As shown in the Tab. 3, our MSTAR achieves 95.71% of MAP on the benchmark.
@@ 4.4 Word-level Scene Text Retrieval
In this part, we present comparison experiments on word-level retrieval. The model is trained on $100k$ images randomly sampled from SynthText-900k [gomez2018single] and 5k images from MLT-5K [nayef2019icdar2019] dataset. The evaluation setting keeps the same as the previous method [wang2021scene]. Note that Deepsolo  [ye2023deepsolo], TG-Bridge  [huang2024bridging], and SPTSv2 [liu2023spts] are well fine-tuned on the training set of Total-Text, CTW and IC15 dataset.
[…]
@@ 4.5 Ablation Study
Ablation study on three core components. The overall results are reported in Tab. 6. We begin by directly training the vision encoder, MLP and multi-modal encoder with standard contrastive learning. The results suggest that this baseline performs poorly on the fine-grained recognition capability, especially on the CTR and IC15 datasets which features small text. Then instructions are adopted to prompt the model to encode queries which leads to improvement. Subsequently, the MIM is added, which leads to a significant improvement of 2.19% on STR and 1.35% on CTW. However, a slight performance drop of 0.46% is observed on the MQTR dataset. This may be because of the use of Hungarian matching, which is effective for word-level queries and instance-level alignment, introduces confusion when handling more complex queries.
Lastly, we incorporate PVE which is designed to improve the retrieval performance of the insalient objects such as small and detailed text instances. The results show a substantial improvement on CTR (4.36%) and IC15 (1.98%), validating the effectiveness of PVE in small text scenarios.
@@
TH+$C$ 59.66 85.17 80.17
TH+WS+$C$ 60.13 85.55 81.21
Table 7: Ablation studies on $\sigma$ algorithm. TH denotes ThresHolding, $C$ denotes Connected Components, and WS denotes WaterShed algorithm.
@@ 5 Discussion
Application of MSTAR for text localization. To further validate the effectiveness of MSTAR, we use it for text localization using Grad-CAM [selvaraju2017grad]. As illustrated in Fig.  3, MSTAR can localize different types of queries. For example, MSTAR accurately identifies the target text instance “dream big” on a t-shirt, distinguishing it from the “dream big” of the shorts in Fig.  3 (a). In addition, MSTAR can also accurately localize curved and dense text instances. As Fig.  3 (e) shows, MSTAR successfully identifies the word “copyright” within a document page image. These results demonstrate that MSTAR can accurately localize text instances without box supervision.
Discussion about the pre-trained model. We discuss how the pre-trained model affects performance on the scale of parameters and data. We tested different variants of CLIP [xue2022language], BLIP [li2022blip], BLIP-2 [li2023blip], and SigLIP [zhai2023sigmoid] on the CTR dataset. The results show that despite the massive pre-training, the models struggle to achieve retrieval on small text instances. Details are provided in the appendix.
Limitations. Although our method shows evident improvements in fine-grained scene text retrieval, it still lags behind box-based methods when handling extremely small and dense text instances. This limitation stems from insufficient positional supervision, which is a common limitation to box-free approaches.
@@ 6 Conclusion
This work investigates the fundamental OCR task of scene text retrieval via a novel box-free paradigm. It incorporates PVE to shift visual attention to fine-grained scene text. Our model demonstrates competitive retrieval performance with state-of-the-art box-based methods while significantly reducing annotation cost. Moreover, for the first time, we study the multi-query scene text retrieval enabling broader real-world applications. Experiments show that neither box-based methods nor general cross-modal retrievers can handle such a challenging task effectively, while our method can serve as a strong baseline for this task. This study is expected to inspire future research on weakly supervised pre-training for visual document foundation models.
Acknowledgements
This work was supported by the National Key Research and Development Program of China (Grant 2022YFC3301703) and the NSFC (Grants 62206104, 62225603).
Figure 4: Qualitive analysis of VLMs to process in-salient text instances , which is introduced in Sec. 1, (a) BLIP-ViT-Large-384, (b) SigLIP-ViT-Base-512, and (c) SigLIP-ViT-Large-384.
@@ A.1 Further Observations on VLMs
To validate the observations that VLMs tend to overlook detailed and in-salient text instances, we conducted additional experiments with more models. For the BLIP, we calculate the ITM score which calculates the one-to-one matching probability between the image and the query. For the SigLIP models, we calculate the dot products between the vision and text embeddings without normalization following the official code. As illustrated in Fig. 4 (a), the BLIP can easily capture the salient text like “dead lakes park” but cannot find the text “old florida”. However, it can capture the “old florida” when the salient regions are covered. Similar observations are presented in Fig. 4 (b) (c).
@@ A.2 Performance of VLMs on Small Text Dataset
To test the pre-trained ability of VLMs on fine-grained perception, we evaluate representative VLMs (CLIP [xue2022language], BLIP [li2022blip], BLIP-2 [li2023blip], SigLIP [zhai2023sigmoid]) on the CTR [veit2016coco] dataset. The CTR dataset is collected from the COCO dataset and features small scene text instances. As the results in Tab.  9, all of the tested VLMs struggle to retrieve images accurately. For instance, the CLIP variants obtain MAP of 8.1% while the BLIP gets 6.9% of MAP. Moreover, increasing model parameters and seen data does not lead to significant performance gains. Among these models, BLIP-2 has the highest MAP of 13.3% of the tested models, indicating limited capacity of VLMs for small text retrieval.
Model Parameters Pre-training Data MAP%
@@
BLIP-2-ViT-Large 452M 129M images 13.3
@@ B.1 Construction details of the MQTR dataset
Word Subset. The word subset includes 5000 images and 200 word queries. The images are sourced from the test set of SVT, CTW, IC15, Total-Text and the CTR. We extract the word-level annotations from the datasets and filter out the words with less than 3 characters (e.g. “st”). After that, 200 word queries are selected according to the word frequency.
Phrase Subset. The images of phrase subset include 1k images from PSTR [zeng2024focus], 1k images manually collected from the Web. Then we use images from HierText [long2023icdar]. The line-level annotations are used and the lines with only one word are filtered out. Lastly, 200 phrase queries are selected according to frequency.
Combined Subset We first use all images collected from the CTR and HierText dataset. Given the word and line annotations of an image, we then filter out the queries less than 3 characters. Then we implement an DFS algorithm to compute all the combinations that contain 2-4 words. Top 200 text combinations on the dataset are selected according to frequency and repeat ratios. Then images paired to these queries are first selected as positive samples. Then images containing words similar to the 200 combined queries are also selected as negative samples according to the edit distance. Lastly, there are 5000 images for the positive samples and negative samples in total.
Semantic Subset The semantic subset is manually collected from the web. we first brainstorm common-used scene text queries and then search for candidate images from the web. In total, the subset consists of 25 queries and 1000 images.
During the selection of multi-word queries (i.e., phrase and combined types), we manually filter out redundant entries. Specifically, queries that are overly repeated or semantically similar are removed to enhance diversity in the final set.